# Sorbonne Université, CNRS, 4 Place Jussieu, 75005 Paris, France

### v1

@@ V The scale of inflation, large field excursions and the tower of light states
It is known that Starobinsky inflation is of large field type implying a breakdown of the effective field theory at a scale below the Planck mass. In particular, as we mentioned earlier, the inflaton value at horizon exit $\phi_{*}$ is about 5 in Planck units, implying the existence of a tower of light states according to the swampland distance conjecture Ooguri:2006in . In order to identify the origin and nature of the tower, it is sufficient to focus on the Kähler potential and its emergence within the string effective action, since the superpotential is suppressed by the inflation scale proportional to $M$, which is constrained by observations to be at least 5 orders of magnitude below the Planck scale. By inspection of its form (18), inflation takes place near the boundary of the Kähler domain $|y|\to 1$, equivalent to large $T_{\rm R}$.
@@
We recall that the $y$- and $z$-dependent part of the Kähler potential corresponds to fields associated with one of the three planes of $T^{6}$ which takes the factorized form $T^{2}\times T^{2}\times T^{2}$, where all fields are at the free fermionic point. Following the derivation of the $N=1$ effective supergravity of four-dimensional superstrings constructed within the free fermionic formulation and translated in a basis of a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ orbifold Antoniadis:1987zk ; Ferrara:1987tp , $z$ is an untwisted field from the third plane while $y$ is twisted in the first two planes Antoniadis:2020txn . Because of the symmetry between all fields defined around the fermionic point, the boundary of the Kähler domain corresponds by the field redefinition $y\to T$ in (47) to a decompactification limit where the area of the third $T^{2}$ $(T_{\rm R})$ becomes large. Indeed $T=\sqrt{G}+ib$ in string units, where $G$ is the determinant of the two-dimensional metric of the third plane and $b$ is the corresponding 2-index antisymmetric tensor which in 2 dimensions has one element. At the fermionic point, the complex structure is unity, corresponding to a square torus of radius $R$, and thus $\sqrt{G}=R^{2}$.
@@
$-2{\partial T\partial\bar{T}\over(T+\bar{T})^{2}}=-2\left({\partial R\over R}\right)^{2}-{1\over 2}{(\partial b)^{2}\over R^{2}}\,.$ (49)
@@
$RM_{p}={1\over g_{s}}(RM_{*})^{2+d/2}\,.$ (52)
As a result, the compactification scale associated to the tower of ‘light’ KK states is $R^{-1}\sim g_{s}e^{-7.5}M_{p}$ which is about four orders of magnitude below the Planck mass, around $10^{14}$ GeV. This is well above the scale of inflation (by an order of magnitude) and does not have any effect in the effective field theory of inflation. Note however that a value of $\phi\simeq 6$ (corresponding to $T_{R}\simeq 400$) would bring the compactification scale of order of the inflation scale invalidating the effective field theory. We would like to emphasize that starting with the Planck determination of $n_{s}$ and noting that in this class of models, $n_{s}\approx 1-2/N_{*}$, where
[…]
@@
$-3{\partial T\partial\bar{T}\over(T+\bar{T})^{2}}=-12\left({\partial R\over R}\right)^{2}-{3\over 4}{(\partial b)^{2}\over R^{4}}\,,$ (55)
where now $T=R^{4}+ib$ in string units. Agreement between (55) with (50) now implies that $d=6$, consistent with the argument above. It follows that the canonically normalized inflaton field is now $\phi=2\sqrt{6}\ln R$, while the compactification scale of six extra dimensions is found using (55) for $d=6$ to be $R^{-1}\sim g_{s}e^{-5}M_{p}$ which is about three orders of magnitude below the Planck mass, around $10^{15}$ GeV. In this case, the effective field theory of inflation breaks around $\phi\sim 10$ corresponding to $T_{R}\simeq 10^{4}$.
@@ VI Conclusions
As noted several times, in addition to a (complex)-inflaton, an additional complex scalar is required. This can be seen from the supersymmetrization of the $R+\alpha R^{2}$ theory Cecotti:1987sa , or from the requirement of achieving inflation within the framework of no-scale supergravity Ellis:2013nxa . All through inflation, the additional directions must sit in stable minima. If the inflaton is associated with the real part of $T$, $T_{\rm R}$, then $T_{\rm I}=0$ is automatically fixed. However, the additional field, $C$, (associated with the sgoldstino) typically has $m_{C}^{2}<0$ and is in fact unstable. We have derived here general conditions for which the dilaton can be used to stabilize the sgoldstino without any additional $C$-dependent corrections to either the Kähler potential or the superpotential. We have also shown that in this construction, the necessary field excursions, $\phi\sim 5$ and $T\sim 40$ are small enough so as to evade the problems which relegate large field inflation models to the swampland, thus turning the swampland into a mirage.

### v2

@@ V The scale of inflation, large field excursions and the tower of light states
It is known that Starobinsky inflation is of large field type implying a breakdown of the effective field theory at a scale below the Planck mass. In particular, as we mentioned earlier, the inflaton value at horizon exit $\phi_{*}$ is about 5 in Planck units, implying the existence of a tower of light states according to the swampland distance conjecture Ooguri:2006in . In order to identify the origin and nature of the tower, it is sufficient to focus on the Kähler potential and its emergence within the string effective action, since the superpotential is suppressed by the inflation scale proportional to $M$, which is fixed by CMB observations using Eq. (2) with $m_{\phi}\propto M$ to be about 5 orders of magnitude below the Planck scale. By inspection of its form (18), inflation takes place near the boundary of the Kähler domain $|y|\to 1$, equivalent to large $T_{\rm R}$.
@@
We recall that the $y$- and $z$-dependent part of the Kähler potential corresponds to fields associated with one of the three planes of $T^{6}$ which takes the factorized form $T^{2}\times T^{2}\times T^{2}$, where all fields are at the free fermionic point. Following the derivation of the $N=1$ effective supergravity of four-dimensional superstrings constructed within the free fermionic formulation and translated in a basis of a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ orbifold Antoniadis:1987zk ; Ferrara:1987tp , $z$ is an untwisted field from the third plane while $y$ is twisted in the first two planes Antoniadis:2020txn . Because of the symmetry between all fields defined around the fermionic point, the boundary of the Kähler domain corresponds by the field redefinition $y\to T$ in (47) to a decompactification limit where the area of the third $T^{2}$ $(T_{\rm R})$ becomes large. Indeed $2T=\sqrt{G}+ib$ in string units, where $G$ is the determinant of the two-dimensional metric of the third plane and $b$ is the corresponding 2-index antisymmetric tensor which in 2 dimensions has one element. At the fermionic point, the complex structure is unity, corresponding to a square torus of radius $R$, and thus $\sqrt{G}=R^{2}$.
@@
$2{\partial T\partial\bar{T}\over(T+\bar{T})^{2}}=-2\left({\partial R\over R}\right)^{2}-{1\over 2}{(\partial b)^{2}\over R^{2}}\,.$ (49)
@@
$RM_{p}={1\over g_{s}}(RM_{*})^{1+d/2}\,.$ (52)
We recall that, in the absence of the dilaton $(g_{s}=1)$, the species scale associated to the KK tower of $d$ flat extra dimensions is of the order of the $(4+d)$-dimensional Planck scale $M_{*}$ and the number of species is given by $N=(RM_{*})^{d}$, so that (51) is reduced to the well known relation $M_{*}=M_{p}/\sqrt{N}$ Dvali:2007hz ; Dvali:2007wp . On the other hand for a perturbative string tower alone $(RM_{*}=1)$, the species scale is the string scale (still denoted $M_{*}$) and the number of species is given by $N=1/g_{s}^{2}$ Dvali:2009ks ; Dvali:2010vm . In the presence of the dilaton with a (modestly) perturbative string coupling, the string scale is slightly below the higher-dimensional Planck scale by a factor of $g_{s}^{1/d}$.
As a result, the compactification scale associated to the tower of ‘light’ KK states is $R^{-1}\sim g_{s}e^{-5}M_{p}$ which is about three orders of magnitude below the Planck mass, around $3\times 10^{15}$ GeV. This is well above the scale of inflation (by two orders of magnitude) and does not have any effect in the effective field theory of inflation. Note however that a value of $\phi\simeq 10$ (corresponding to $T_{R}\simeq 10^{4}$) would bring the compactification scale near the inflation scale ($H\sim M/\sqrt{12}\sim 5\times 10^{-6}M_{P}$) and larger values of $\phi$ would invalidate the effective field theory. We would like to emphasize that starting with the Planck determination of $n_{s}$ and noting that in this class of models, $n_{s}\approx 1-2/N_{*}$, where
$N_{*}\simeq-\int^{\phi_{*}}_{\phi_{\rm end}}\frac{1}{\sqrt{2\epsilon}}d\phi$ (53)
@@
Note that it has already been pointed out in the literature that modest super-Planckian displacements are not problematic for the validity of the effective field theory, as long as the species scale remains above the Hubble scale Scalisi:2018eaz ; vandeHeisteeg:2023uxj . Here we impose a stronger condition that the scale of the KK tower (compactification scale) should remain above the Hubble scale during inflation.
@@
$3{\partial T\partial\bar{T}\over(T+\bar{T})^{2}}=-12\left({\partial R\over R}\right)^{2}-{3\over 4}{(\partial b)^{2}\over R^{4}}\,$, (55)
where now $2T=R^{4}+ib$ in string units. Agreement between (55) with (50) now implies that $d=6$, consistent with the argument above. It follows that the canonically normalized inflaton field is now $\phi=2\sqrt{6}\ln R$, while the compactification scale of six extra dimensions is found using (55) for $d=6$ to be $R^{-1}\sim g_{s}e^{-5\sqrt{2/3}}M_{p}$ which is less than three orders of magnitude below the Planck mass, around $8\times 10^{15}$ GeV. In this case, the effective field theory of inflation breaks around $\phi\sim 12$ again corresponding to $T_{R}\simeq 10^{4}$.
We conclude this section by deriving a relation between the number of e-folds $N_{e}$ during inflation and the number of species $N$, as advocated in the introduction. $N_{e}$ is directly related to the initial value of the inflaton through Eq. (53) by replacing $\phi_{\star}$ with the initial inflaton field value at the beginning of inflation, $\phi_{0}$, and by replacing $N_{\star}$ with $N_{e}$. More precisely, in the examples studied above with scalar potential (1) in the Starobinsky model and (21) for the string model, on obtain (assuming $\phi_{\rm end}\approx 0)$:
[…]
On the other hand $\phi_{0}$ determines the number of species $N$ as follows (assuming for simplicity $\phi_{\text{end}}\approx 0$). Since $\phi_{0}>\phi_{\star}$ is large in Planck units, the KK tower opens up at a lower scale than estimated above, given by (52) with $\phi_{\star}$ replaced by $\phi_{0}$. The species scale is of the order of the higher dimensional Planck scale $M_{*}$ and the number of species is given by $N=(RM_{*})^{d}$, where $d$ is the number of extra dimensions in the KK tower. From Eq. (50) one obtains that $\phi=[d(d+2)/2]\ln(RM_{*})$ and thus
$N=e^{2\phi_{0}/(d+2)}.$ (57)
Combining (56) with (57), one obtains a relation between $N_{e}$ and $N$, of the form $N_{e}\sim N^{b}$ where the exponent $b$ is of order unity and $b\gtrsim 1$, at least for the models considered here.
@@ VI Conclusions
As noted several times, in addition to a (complex)-inflaton, an additional complex scalar is required. This can be seen from the supersymmetrization of the $R+\alpha R^{2}$ theory Cecotti:1987sa , or from the requirement of achieving inflation within the framework of no-scale supergravity Ellis:2013nxa . All through inflation, the additional directions must sit in stable minima. If the inflaton is associated with the real part of $T$, $T_{\rm R}$, then $T_{\rm I}=0$ is automatically fixed. However, the additional field, $C$, (associated with the sgoldstino) typically has $m_{C}^{2}<0$ and is in fact unstable. We have derived here general conditions for which the dilaton can be used to stabilize the sgoldstino without any additional $C$-dependent corrections to either the Kähler potential or the superpotential. We have also shown that in this construction, the necessary field excursions, $\phi\sim 5$ and $T\sim 75$ are small enough so as to evade the problems which relegate large field inflation models to the swampland, thus turning the swampland into a mirage.